# Fractionalization as an alternate to charge ordering in electronic insulators

### v1

@@ Appendix D Results about non-Abelian TOs
We begin in App. D.1 by arguing a result about braiding with translation invariant anyons in a SET $\mathcal{C}$ which is assumed to be symmetric under the full symmetry group $G$ of Eq. (4). We will see in the special case that translation does not permute anyons $a=\mathcal{T}a$, where $a$ is the background anyon. This result will also later prove useful when we consider the uniqueness of $\mathbb{Z}_{2q}$ gauge theory even if translation is allowed to permute anyons.
We then switch gears to prove some results bounding the numbers of anyons in a non-Abelian theory. In App. D.2 we will first use the results of App. C to prove a few things about the size of the subgroup of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ when $\mathcal{C}$ is a time reversal invariant TO. Since $|\mathcal{C}|\geq|\mathcal{A}|$ with equality only when $\mathcal{C}$ is an Abelian TO this gives us a powerful tool to bound the size of $|\mathcal{C}|$. Finally in App. D.3 we will use the results of App. D.2 to demonstrate that every minimal order which satisfies the full symmetry group $G$ of Eq. (4) at filling $\nu=1/q$ with $q$ even must contain exactly $8q^{2}$ anyons.
@@ D.1 SETs of $G$
In this section we will offer an argument for the following result. Suppose that $\mathcal{C}$ is a SET of the full group $G$ in Eq. (4). Then the background anyon and its time reversal partner must braid identically with all translation invariant anyons, i.e.
@@
The argument for Eq. (46) is very straightforward, and relies on the fact that our state must be invariant under time reversal. Suppose that we think of our state as possessing a background anyon $a$ at the center of every plaquette. Then under time reversal the state will be transformed to a state with $\mathcal{T}a$ at the center of every plaquette. Since we have not broken time reversal symmetry we should not be able to detect the difference between these two states. As the way we detect differences between anyons is via braiding, naïvely we should expect that this means $a$ and $\mathcal{T}a$ must braid identically with all anyons. However, there is a subtlety to consider if translation is allowed to permute anyons.
Let us then be careful about defining the braiding of an anyon with the background anyon. First we create a pair of anyons $d$ and $\overline{d}$. Now we want to take $d$ around a plaquette, thus braiding it with $a$, but leave $\overline{d}$ in place. This will require first defining a string operator which is able to translate only the anyon $d$ by one lattice site in either direction. Let us specialize to the $x$-direction and suppose such an operator exists. The existence of such a local operator will imply that $d$ and $d$ translated by one lattice site, call this $T_{x}d$, must be in the same superselection sector. But, this is precisely the definition of anyon type Kitaev (2006), so the existence of such an operator implies that $d=T_{x}d$. The same will be true in the $y$-direction. Thus we see that our ability to take $d$ around a plaquette and leave $\overline{d}$ in place is exactly equivalent to the statement that $T_{x}d=d=T_{y}d$.
For an anyon $d$ which is left invariant by translation this process is possible, and we can find a string operator which takes $d$ around a plaquette. This will lead to the wavefunction acquiring a phase of $\theta_{a,d}$. Doing the same thing in the time reversed state will lead to a phase of $\theta_{\mathcal{T}a,d}$. Since we have not broken time reversal symmetry, these two phases should be identical and we have the result. This is illustrated in Fig. LABEL:fig:background.
@@ D.2 Bounds on time reversal invariant TOs
In this subsection we will focus on proving results about a generic time reversal invariant TO; we assume no other symmetries.
@@
We note that when restricted to $\mathcal{A}\subseteq\mathcal{C}$ the braiding phase $\theta\colon\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{S}^{1}$ is still a symmetric group homomorphism. However, we no longer know that braiding is non-degenerate, even up to the fundamental fermion. Indeed it may be the case that there is some $a\in\mathcal{A}$ which braids trivially with all other anyons in $\mathcal{A}$, but nontrivially with some non-Abelian anyon in $\mathcal{C}\setminus\mathcal{A}$. Having noted that we make the following definition.
@@
The set $\mathcal{N}\subseteq\mathcal{A}$ forms a subgroup of $\mathcal{A}$. Further, all elements of $\mathcal{N}$ are bosons or fermions. If we consider $\mathcal{A}/\mathcal{N}$ as a group, then braiding is non-degenerate on this group and $|\mathcal{A}|=|\mathcal{A}/\mathcal{N}||\mathcal{N}|$.
@@
Since $\mathcal{N}$ is precisely the subgroup of $\mathcal{A}$ that braids trivially with every element in $\mathcal{A}$ we see that braiding must be non-degenerate on $\mathcal{A}/\mathcal{N}$.
@@
Now suppose that $a\in\mathcal{A}$, then $d_{a}=1$. Since $\mathcal{T}$ is a symmetry of $\mathcal{C}$ we must have that $d_{\mathcal{T}a}=d_{a}=1$ Barkeshli et al. (2019). Thus we see that $\mathcal{T}(\mathcal{A})\subseteq\mathcal{A}$. Since we take $\mathcal{T}^{2}$ to be the identity we must then have that $\mathcal{T}(\mathcal{A})=\mathcal{A}$. Thus $\mathcal{T}$ restricts to a bijection on $\mathcal{A}$. Further, since $\mathcal{T}$ is a symmetry of $\mathcal{C}$ we must also have that the fusion data of $\mathcal{C}$ is left invariant by action of $\mathcal{T}$ Barkeshli et al. (2019). Thus we must also have that $\mathcal{T}(ab)=(\mathcal{T}a)(\mathcal{T}b)$ for $a,b\in\mathcal{A}$. Since $\mathcal{T}\colon\mathcal{A}\rightarrow\mathcal{A}$ is a bijective homomorphism it must restrict to an automorphism on $\mathcal{A}$. ∎
@@
We further know that $\mathcal{A}/\mathcal{N}$ is an Abelian group where braiding is non-degenerate. Then we can run the same arguments of App. C to show that if there is some element of $\mathcal{A}/\mathcal{N}$ that has order $n$ then $|\mathcal{A}/\mathcal{N}|=(kn)^{2}/2$ or $(nk)^{2}$ for some $k$. Further if there are no anyons $a\in\mathcal{A}$ such that $\mathcal{T}a=af$ for some fermion $f\in\mathcal{N}$ then we know that every time-reversal invariant anyon in $\mathcal{A}/\mathcal{N}$ must have $0=2\theta_{a}=\theta_{a,a}$. Thus the arguments of Lem. C.20 reveal that $|\mathcal{A}/\mathcal{N}|$ must be a square and hence $|\mathcal{A}/\mathcal{N}|=(nk)^{2}$. ∎
@@ D.3 Proof by contradiction when $q$ even
In this appendix we will show that $|\mathcal{A}|=8q^{2}$ for any putative minimal order when $q$ is even. In the main text we assumed that $\mathcal{T}a=a$ in order to show this; here we do not assume this.
Suppose that $\mathcal{C}\in mG_{\rm GSD}$ is the minimal TO for the group $G$ in Eq. (4). Then let $\mathcal{A}\subseteq\mathcal{C}$ be the group of Abelian anyons and $\mathcal{N}\subseteq\mathcal{A}$ be the subgroup of $\mathcal{A}$ containing all anyons that braid trivially with every anyon in $\mathcal{A}$. Let $c$ be the fundamental fermion.
@@
First, it is evident that the order of $v$ in $\mathcal{A}/\mathcal{N}$ is a multiple of $q$ because $\theta_{v,a}=2\pi/q$. Suppose the order of $v$ in $\mathcal{A}/\mathcal{N}$ is $k_{v}q$ with $k_{v}>1$. Then $|\mathcal{A}/\mathcal{N}|$ is at least $k_{v}^{2}q^{2}$, for the following reason: without loss of generality, we can assume there is an Abelian anyon $x$ with $\theta_{x,v}=2\pi/k_{v}q$. Consider the set $\mathcal{S}=\{x^{k}v^{l}|0\leq k,l\leq k_{v}q-1\}$. For any $l$, $x^{k}v^{l}$ braids nontrivially with $v$ as long as $k\neq 0$. If $k=0$, then $v^{l}$ braids nontrivially with $x$ if $l\neq 0$. Thus the braiding on $\mathcal{S}$ is non-degenerate, and $\mathcal{S}$ must be fully contained in $\mathcal{A}/\mathcal{N}$. Because $|\mathcal{N}|$ is at least 2 (it contains $\{1,c\}$), we find $|\mathcal{A}|\geq 2k_{v}^{2}q^{2}>8q^{2}$ if $k_{v}>2$. Thus from the minimality assumption, we conclude $k_{v}=1,2$. If $k_{v}=2$, then the minimality assumption requires $|\mathcal{A}|=8q^{2}$ and $\mathcal{A}=\mathcal{S}\boxtimes\{1,c\}$, which is entirely Abelian. This is the first case in the statement. Thus in the following we set $k_{v}=1$.
@@
Next we assume $\mathcal{N}\neq\{1,c\}$, which implies $|\mathcal{N}|\geq 4$. If $a^{q}\in\mathcal{A}/\mathcal{N}$, then $|\mathcal{A}/\mathcal{N}|\geq 2q^{2}$ (since $\mathcal{A}/\mathcal{N}$ already contains $\mathcal{S}$). Together $|\mathcal{A}|\geq 2q^{2}\times 4=8q^{2}$. However, in this case there has to be at least one (non-Abelian) anyon in $\mathcal{C}$ to braid nontrivially with $\mathcal{N}/\{1,c\}$, so $|\mathcal{C}|$ must be strictly greater than $8q^{2}$, violating the minimality assumption. We thus conclude that $a^{q}\in\mathcal{N}$.
@@
Because $a^{q}$ has charge $1\pmod{2}$, the order of $a^{q}\in\mathcal{N}\subseteq\mathcal{A}$ must be a multiple of two, i.e. $n_{a}=2qr$ for some integer $r$. Further, we find
@@
If $|\mathcal{A}/\mathcal{N}|=q^{2}$, then for every $\sigma\in\mathcal{C}\setminus\mathcal{A}$ we must have $\theta_{a^{q},\sigma}=\pi$. Further, for every pair of non-Abelian anyons $\sigma,\sigma^{\prime}\in\mathcal{C}\setminus\mathcal{A}$ we must have that
@@
for some $b\in\mathcal{A}$, where we have included fusion multiplicities. But we must have that $N^{b}_{\sigma\sigma^{\prime}}=N^{\sigma^{\prime}}_{\overline{\sigma}b}$ (see Kitaev (2006)). We further know that the fusion multiplicities for $\overline{\sigma}\times b$ are always zero or one since $b$ is an Abelian anyon, and thus $N^{\sigma^{\prime}}_{\overline{\sigma}b}=0,1$. This tells us that the above are either zero or one. Using Eq. (21) of Barkeshli et al. (2019) we further see that this means:
@@
Now we invoke a theorem in Barkeshli et al. (2019) (see Eq. (244)), that because of the $\mathbb{Z}_{2}$ grading on $\mathcal{C}$, we must have $\mathcal{D}_{\mathcal{C}_{0}}^{2}=\mathcal{D}_{\mathcal{C}_{\pi}}^{2}=4q^{2}$. Thus there are precisely $2q^{2}$ anyons in $\mathcal{C}_{\pi}$. For any $\sigma\in\mathcal{C}_{\pi}$, all anyons of the form $\mathcal{A}\times\sigma$ are still in $\mathcal{C}_{\pi}$, so to get $2q^{2}$ anyons instead of $4q^{2}$, there must be precisely one nontrivial anyon $x\in\mathcal{A}$ such that $x\times\sigma=\sigma$, and we have shown earlier that charge conservation implies $x$ must be $a^{q}c$. We thus conclude that $\sigma\times a^{q}c=\sigma$. ∎
@@
So without loss of generality we will take $\gamma$ to have charge zero modulo one, where $\theta_{a,\gamma}=\pi p/q$ for some odd $p=2p^{\prime}+1$. We then consider $v^{k}\times\gamma$, we see that:
@@
Next, we prove $\theta_{\gamma}=0$. For this we will need the following relation:
@@
which was shown in Lapa and Levin (2019) for any fermionic TO using arguments about gauging the fermion parity. Since we have assumed time reversal invariance the left hand side must equal one.
@@
We will now do some algebra, building off of our previous results. First we have:
@@
The sums over $l$ will be zero unless $k$ is a multiple of $q$, so $k$ can only take two values $k=0,q$ and the sum over $l$ gives $q$. Note that $q^{2}\theta_{a}=\theta_{a^{q}}=1$. With $Q_{a}=1/q$,
@@
Now consider the fusion product of $\gamma\times\gamma$, we know that it is equal to $b+ba^{q}c$ for some $b\in\mathcal{A}$. Further since $\theta_{a,\gamma}=\pi Q_{a}$, we must have that $b$ braids identically to the vison, which implies $b=v$ or $b=vc$ by braiding non-degeneracy. In addition, $Q_{\gamma}=0$ mod 2 implies $Q_{v}=0$ mod 2, which shows $b=v$. ∎
@@
We can see this by condensing the bosonic vison $v$, forming a child TO $\mathcal{C}^{\prime}$. This will confine any anyon which braids nontrivially with $v$, i.e. has charge not equal to an integer. It will further identify all powers of $v$ with the vacuum. It is then clear that:
@@
and $(a^{q}c)^{2}=1$ so $\mathcal{C}^{\prime\prime}$ is closed under fusion. We note that $\mathcal{C}^{\prime\prime}$ has the fusion rules of the Ising TO, but the Ising anyon $\gamma$ has $\theta_{\gamma}=0$. However, all TOs with these fusion rules were classified in Kitaev (2006), and in all of them the non-Abelian Ising anyon satisfies $8\theta_{\gamma}=\pi\pmod{2\pi}$.

### v2

@@ Appendix D Results about non-Abelian TOs
We begin in Appendix D.1 by arguing a result about braiding with translation invariant anyons in a SET $\mathcal{C}$ which is assumed to be symmetric under the full symmetry group $G$ of Eq. (4). We see in the special case that translation does not permute anyons $a=\mathcal{T}a$, where $a$ is the background anyon. This result will also later prove useful when we consider the uniqueness of $\mathbb{Z}_{2q}$ gauge theory even if translation is allowed to permute anyons.
We then switch gears to prove some results bounding the numbers of anyons in a non-Abelian theory. In Appendix D.2 we first use the results of Appendix C to prove a few things about the size of the subgroup of Abelian anyons $\mathcal{A}\subseteq\mathcal{C}$ when $\mathcal{C}$ is a time reversal invariant TO. Since $|\mathcal{C}|\geq|\mathcal{A}|$ with equality only when $\mathcal{C}$ is an Abelian TO this gives us a powerful tool to bound the size of $|\mathcal{C}|$. Finally in Appendix D.3 we use the results of Appendix D.2 to demonstrate that every minimal order which satisfies the full symmetry group $G$ of Eq. (4) at filling $\nu=1/q$ with $q$ even must contain exactly $8q^{2}$ anyons.
@@ D.1 SETs of $G$
In this section we offer an argument for the following result. Suppose that $\mathcal{C}$ is a SET of the full group $G$ in Eq. (4). Then the background anyon and its time-reversal partner must braid identically with all translation invariant anyons, i.e.
@@
The argument for Eq. (46) is very straightforward, and relies on the fact that our state must be invariant under time reversal. Suppose that we think of our state as possessing a background anyon $a$ at the center of every plaquette. Then under time reversal the state will be transformed to a state with $\mathcal{T}a$ at the center of every plaquette. Since we have not broken time-reversal symmetry we should not be able to detect the difference between these two states. As the way we detect differences between anyons is via braiding, naïvely we should expect that this means $a$ and $\mathcal{T}a$ must braid identically with all anyons. However, there is a subtlety to consider if translation is allowed to permute anyons.
Let us then be careful about defining the braiding of an anyon with the background anyon. First we create a pair of anyons $d$ and $\overline{d}$. Now we want to take $d$ around a plaquette, thus braiding it with $a$, but leave $\overline{d}$ in place. This will require first defining a string operator which is able to translate only the anyon $d$ by one lattice site in either direction. Let us specialize to the $x$ direction and suppose such an operator exists. The existence of such a local operator will imply that $d$ and $d$ translated by one lattice site, call this $T_{x}d$, must be in the same superselection sector. But, this is precisely the definition of anyon type Kitaev (2006), so the existence of such an operator implies that $d=T_{x}d$. The same will be true in the $y$ direction. Thus we see that our ability to take $d$ around a plaquette and leave $\overline{d}$ in place is exactly equivalent to the statement that $T_{x}d=d=T_{y}d$.
For an anyon $d$ which is left invariant by translation this process is possible, and we can find a string operator which takes $d$ around a plaquette. This will lead to the wavefunction acquiring a phase of $\theta_{a,d}$. Doing the same thing in the time reversed state will lead to a phase of $\theta_{\mathcal{T}a,d}$. Since we have not broken time-reversal symmetry, these two phases should be identical and we have the result. This is illustrated in Fig. LABEL:fig:background.
@@ D.2 Bounds on time reversal invariant TOs
In this section we will focus on proving results about a generic time reversal invariant TO; we assume no other symmetries.
@@
We note that when restricted to $\mathcal{A}\subseteq\mathcal{C}$ the braiding phase $\theta\colon\mathcal{A}\times\mathcal{A}\rightarrow\mathbb{S}^{1}$ is still a symmetric group homomorphism. However, we no longer know that braiding is nondegenerate, even up to the fundamental fermion. Indeed it may be the case that there is some $a\in\mathcal{A}$ which braids trivially with all other anyons in $\mathcal{A}$, but nontrivially with some non-Abelian anyon in $\mathcal{C}\setminus\mathcal{A}$. Having noted that we make the following definition.
@@
The set $\mathcal{N}\subseteq\mathcal{A}$ forms a subgroup of $\mathcal{A}$. Furthermore, all elements of $\mathcal{N}$ are bosons or fermions. If we consider $\mathcal{A}/\mathcal{N}$ as a group, then braiding is nondegenerate on this group and $|\mathcal{A}|=|\mathcal{A}/\mathcal{N}||\mathcal{N}|$.
@@
Since $\mathcal{N}$ is precisely the subgroup of $\mathcal{A}$ that braids trivially with every element in $\mathcal{A}$ we see that braiding must be nondegenerate on $\mathcal{A}/\mathcal{N}$.
@@
Now suppose that $a\in\mathcal{A}$, then $d_{a}=1$. Since $\mathcal{T}$ is a symmetry of $\mathcal{C}$ we must have that $d_{\mathcal{T}a}=d_{a}=1$ Barkeshli et al. (2019). Thus we see that $\mathcal{T}(\mathcal{A})\subseteq\mathcal{A}$. Since we take $\mathcal{T}^{2}$ to be the identity we must then have that $\mathcal{T}(\mathcal{A})=\mathcal{A}$. Thus $\mathcal{T}$ restricts to a bijection on $\mathcal{A}$. Furthermore, since $\mathcal{T}$ is a symmetry of $\mathcal{C}$ we must also have that the fusion data of $\mathcal{C}$ is left invariant by action of $\mathcal{T}$ Barkeshli et al. (2019). Thus we must also have that $\mathcal{T}(ab)=(\mathcal{T}a)(\mathcal{T}b)$ for $a,b\in\mathcal{A}$. Since $\mathcal{T}\colon\mathcal{A}\rightarrow\mathcal{A}$ is a bijective homomorphism it must restrict to an automorphism on $\mathcal{A}$. ∎
@@
We further know that $\mathcal{A}/\mathcal{N}$ is an Abelian group where braiding is nondegenerate. Then we can run the same arguments of Appendix C to show that if there is some element of $\mathcal{A}/\mathcal{N}$ that has order $n$ then $|\mathcal{A}/\mathcal{N}|=(kn)^{2}/2$ or $(nk)^{2}$ for some $k$. Further if there are no anyons $a\in\mathcal{A}$ such that $\mathcal{T}a=af$ for some fermion $f\in\mathcal{N}$ then we know that every time-reversal invariant anyon in $\mathcal{A}/\mathcal{N}$ must have $0=2\theta_{a}=\theta_{a,a}$. Thus the arguments of Lem. C.20 reveal that $|\mathcal{A}/\mathcal{N}|$ must be a square and hence $|\mathcal{A}/\mathcal{N}|=(nk)^{2}$. ∎
@@ D.3 Proof by contradiction when $q$ even
In this appendix we show that $|\mathcal{A}|=8q^{2}$ for any putative minimal order when $q$ is even. In the main text we assumed that $\mathcal{T}a=a$ in order to show this; here we do not assume this.
Suppose that $\mathcal{C}\in mG^{(\nu)}_{\rm GSD}$ is the minimal TO for the group $G$ in Eq. (4). Then let $\mathcal{A}\subseteq\mathcal{C}$ be the group of Abelian anyons and $\mathcal{N}\subseteq\mathcal{A}$ be the subgroup of $\mathcal{A}$ containing all anyons that braid trivially with every anyon in $\mathcal{A}$. Let $c$ be the fundamental fermion.
@@
First, it is evident that the order of $v$ in $\mathcal{A}/\mathcal{N}$ is a multiple of $q$ because $\theta_{v,a}=2\pi/q$. Suppose the order of $v$ in $\mathcal{A}/\mathcal{N}$ is $k_{v}q$ with $k_{v}>1$. Then $|\mathcal{A}/\mathcal{N}|$ is at least $k_{v}^{2}q^{2}$, for the following reason: without loss of generality, we can assume there is an Abelian anyon $x$ with $\theta_{x,v}=2\pi/k_{v}q$. Consider the set $\mathcal{S}=\{x^{k}v^{l}|0\leq k,l\leq k_{v}q-1\}$. For any $l$, $x^{k}v^{l}$ braids nontrivially with $v$ as long as $k\neq 0$. If $k=0$, then $v^{l}$ braids nontrivially with $x$ if $l\neq 0$. Thus the braiding on $\mathcal{S}$ is nondegenerate, and $\mathcal{S}$ must be fully contained in $\mathcal{A}/\mathcal{N}$. Because $|\mathcal{N}|$ is at least 2 (it contains $\{1,c\}$), we find $|\mathcal{A}|\geq 2k_{v}^{2}q^{2}>8q^{2}$ if $k_{v}>2$. Thus from the minimality assumption, we conclude $k_{v}=1,2$. If $k_{v}=2$, then the minimality assumption requires $|\mathcal{A}|=8q^{2}$ and $\mathcal{A}=\mathcal{S}\boxtimes\{1,c\}$, which is entirely Abelian. This is the first case in the statement. Thus in the following we set $k_{v}=1$.
@@
Next we assume $\mathcal{N}\neq\{1,c\}$, which implies $|\mathcal{N}|\geq 4$. If $a^{q}\notin\mathcal{N}$, then $|\mathcal{A}/\mathcal{N}|\geq 2q^{2}$ (since $\mathcal{A}/\mathcal{N}$ already contains $\mathcal{S}$). Together $|\mathcal{A}|\geq 2q^{2}\times 4=8q^{2}$. However, in this case there has to be at least one (non-Abelian) anyon in $\mathcal{C}$ to braid nontrivially with $\mathcal{N}/\{1,c\}$, so $|\mathcal{C}|$ must be strictly greater than $8q^{2}$, violating the minimality assumption. We thus conclude that $a^{q}\in\mathcal{N}$.
@@
Because $a^{q}$ has charge $1\pmod{2}$, the order of $a^{q}\in\mathcal{N}\subseteq\mathcal{A}$ must be a multiple of two i.e., $n_{a}=2qr$ for some integer $r$. Furthermore, we find
@@
If $|\mathcal{A}/\mathcal{N}|=q^{2}$, then for every $\sigma\in\mathcal{C}\setminus\mathcal{A}$ we must have $\theta_{a^{q},\sigma}=\pi$. Furthermore, for every pair of non-Abelian anyons $\sigma,\sigma^{\prime}\in\mathcal{C}\setminus\mathcal{A}$ we must have that
@@
for some $b\in\mathcal{A}$, where we have included fusion multiplicities. But we must have that $N^{b}_{\sigma\sigma^{\prime}}=N^{\sigma^{\prime}}_{\overline{\sigma}b}$ (see Ref. Kitaev (2006)). We further know that the fusion multiplicities for $\overline{\sigma}\times b$ are always zero or one since $b$ is an Abelian anyon, and thus $N^{\sigma^{\prime}}_{\overline{\sigma}b}=0,1$. This tells us that the above are either zero or one. Using Eq. (21) of Ref. Barkeshli et al. (2019) we further see that this means:
@@
Now we invoke a theorem in Barkeshli et al. (2019) [see Eq. (244)], that because of the $\mathbb{Z}_{2}$ grading on $\mathcal{C}$, we must have $\mathcal{D}_{\mathcal{C}_{0}}^{2}=\mathcal{D}_{\mathcal{C}_{\pi}}^{2}=4q^{2}$. Thus there are precisely $2q^{2}$ anyons in $\mathcal{C}_{\pi}$. For any $\sigma\in\mathcal{C}_{\pi}$, all anyons of the form $\mathcal{A}\times\sigma$ are still in $\mathcal{C}_{\pi}$, so to get $2q^{2}$ anyons instead of $4q^{2}$, there must be precisely one nontrivial anyon $x\in\mathcal{A}$ such that $x\times\sigma=\sigma$, and we have shown earlier that charge conservation implies $x$ must be $a^{q}c$. We thus conclude that $\sigma\times a^{q}c=\sigma$. ∎
@@
So without loss of generality we take $\gamma$ to have charge zero modulo one, where $\theta_{a,\gamma}=\pi p/q$ for some odd $p=2p^{\prime}+1$. We then consider $v^{k}\times\gamma$, we see that:
@@
Next, we prove $\theta_{\gamma}=0$. For this we need the following relation:
@@
which was shown in Ref. Lapa and Levin (2019) for any fermionic TO using arguments about gauging the fermion parity. Since we have assumed time-reversal invariance the left-hand side must equal one.
@@
We now do some algebra, building off of our previous results. First we have:
@@
The sums over $l$ will be zero unless $k$ is a multiple of $q$, so $k$ can only take two values $k=0,q$ and the sum over $l$ gives $q$. Note that $q^{2}\theta_{a}=\theta_{a^{q}}=0$. With $Q_{a}=1/q$,
@@
Now consider the fusion product of $\gamma\times\gamma$, we know that it is equal to $b+ba^{q}c$ for some $b\in\mathcal{A}=\langle a,v\rangle\times\{1,c\}$. Since $Q_{\gamma}=0\pmod{2}$ this means $\theta_{\gamma,v}=0$. This, along with the fact that $\theta_{\gamma,a}=\pi Q_{a}$, means that any fusion product $b\in\gamma\times\gamma$ must be such that $\theta_{b,v}=0$ and $\theta_{b,a}=2\pi Q_{a}$. Thus $b$ braids identically to the vison, which implies $b=v$ or $b=vc$ by braiding nondegeneracy. In addition, $Q_{\gamma}=0$ mod 2 implies $Q_{v}=0$ mod 2, which shows $b=v$. ∎
@@
We can see this by condensing the bosonic vison $v$, forming a child TO $\mathcal{C}^{\prime}$. This will confine any anyon which braids nontrivially with $v$, i.e., has charge not equal to an integer. It will further identify all powers of $v$ with the vacuum. It is then clear that:
@@
and $(a^{q}c)^{2}=1$ so $\mathcal{C}^{\prime\prime}$ is closed under fusion. We note that $\mathcal{C}^{\prime\prime}$ has the fusion rules of the Ising TO, but the Ising anyon $\gamma$ has $\theta_{\gamma}=0$. However, all TOs with these fusion rules were classified in Ref. Kitaev (2006), and in all of them the non-Abelian Ising anyon satisfies $8\theta_{\gamma}=\pi\pmod{2\pi}$.